OpenStudy (anonymous):

For the function f(x) = 3x + 3, find x such that f(x) = 18 A. 6 B. 57 C. 5 D. 3

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks

OpenStudy (gorv):

3x+3=18 3x=15 x=5

OpenStudy (anonymous):

so c

OpenStudy (gorv):

@imawesome101 C is correct one

OpenStudy (anonymous):

Translate f(x) = 3x + 3 to 18 = 3x + 3 Now subtract 3 from both sides 18-3 = 3x +3-3 15 = 3x Now divide both sides by 3 15/3 = 3x/3 5 = x Replace x with 5 in the original equation 18 = 3(5) + 3 18 = 15 + 3 18 = 18 x = 5

OpenStudy (anonymous):

thanks